If one side of a hexagon is 4cm what is the radius and apothem ?
1
Expert's answer
2012-11-27T09:09:28-0500
360/6=60, so we have regular triasngles made of side andradiuses.
That means, that radius is 4 sm.
Apothem is height of triangle,for regular it is a*cos30=a/2=2 sm